Mathematics 133 – Intensive Calculus for Science 1 Things to Know – Quiz 3 September 20, 2000

The third graded quiz of the semester will be given in class on Friday, September 22. There will be three or four questions similar to problems from the third problem set. You should know the following material.

1) Trigonometric functions. Know:

a) Radian measure for angles and how to determine the signs of sin θ, cos θ, tan θ for an angle θ in radians

b) How to sketch graphs for sinusoidal oscillations

y = A sin(Bx) + C y = A cos(Bx) + C

and the meanings of A, B, C

c) How to find a formula for a sinusoidal oscillation, given the graph. (Note: we have not talked about shifting the period to start at points other than 0, so that will not be included)

2) Polynomials and rational functions. Know:

a) How to determine the shape of the graph of a polynomial for big x from the equation.

b) How to determine equations for polynomials given information from a graph such as x- and y-axis intercepts.

c) How to determine where rational functions have vertical and horizontal asymptotes

Good Review Problems:

1) Section 1.9/problems like 1-9, 18, 20 - 28, 30

2) Section 1.10/problems like 3-6, 9-11, 16, 17